Find the slope of the line that goes through the given points
(-1,1)(1/2,4)

Answer by Cromlix(4381)

Hi there,
Gradient = y2 - y1/x2 - x1
Gradient = 4 - 1/1/2 - (-1)
Gradient = 4 - 1/1/2 + 1
Gradient = 3/3/2
Gradient = 6/3
Gradient = 2
This is the slope.
Hope this helps :-)